OpenStudy (anonymous):

can anyone help me step by step please!! 3(-9n + 5) + -4 = -16

OpenStudy (anonymous):

-5x+20+2=-13 and to check- -5x+22=-13 -5(7-4)+2=-13? -5x=-35 -35+20+2=-13? x=7 -15+2=-13 -13=-13!

OpenStudy (anonymous):

-5(x – 4) + 2 = -13 first you distributed the -5 and you get -5x+20 +2=-13 combine like terms -5x+22=-13 subtract 22 from both sides -5x=-35 divide each side by -5 you get x=7

OpenStudy (anonymous):

@Tahmina i need with like 2 more problems so can you stay and help me!! please!!! :) :) :)

OpenStudy (anonymous):

@pinkberry501 you have to explain each step so he/she gets it okay? :)

OpenStudy (anonymous):

okay ill stay but you have to give me time i have my things to do also :)

OpenStudy (anonymous):

@Tahmina um you don't need to correct what I did. Thanks! :)

OpenStudy (anonymous):

@Tahmina & @pinkberry501 :) thanks for all the help!

OpenStudy (anonymous):

u welcome